half-life of silicon 32 is 710 years....40 grams present..
how much in 200 years?

Answer by khwang(438)

Let Q = Qo e^(-kt),
when t =0, Q = 40, so Qo = 40,
So, Q = 40 e^(-kt)
when t=710, Q = 1/2 Q0 = 20,
solve 20 = 40 e^(-710k), or 1/2 = e^(-710k)
we get k = ln2/710 = 0.000976264 = 9.76 *10^(-4)
hence,Q = 40 e^(- 9.76 *10^(-4)t)
when t = 200 yrs, Q = 40 e^(- 9.76 *10^(-4)*200)
= 32.91 grams
Kenny